# Resolving the Body-Order Paradox of Machine Learning Interatomic Potentials

### v1

@@ V Explicit resolution of reference body-orders
Figure 3 shows that the DFT body-order trends are quickly captured by both MACE and PET as the $m$-mers to the training set. In fact, for both $\rho$ 8-mers, near-complete resolution is achieved for $m\leq 6$ at 0.01 added $m$-mer proportion, and further increase in the proportion quickly resolves the $6\leq m\leq 8$ contributions, faster for PET than MACE. This proves that while the MACE and PET architectures have sufficient flexibility to learn the reference body-order trends, without explicit resolution of the body-orders, their inherent tendencies result in trends that are drastically different from the reference. In SOAP-BPNN, $m\leq 4$ contributions are also resolved quickly at 0.01 added $m$-mer proportion, akin to the other two models. SOAP-BPNN, however, struggles to capture the higher body-order contributions at all added $m$-mer proportions, although it still gradually approaches the reference values. Better resolution is achieved for high $\rho$ 8-mers than the low $\rho$ 8-mers. In all MLIPs, the resolution takes place consistently faster for lower $m$, which must be related to how the dimensionality of the potential energy surfaces of $m$-mers exponentially increases with $m$.
@@
Figure 4 shows the energy and force RMSEs for the 8-mers and their $m$-mers. The RMSEs for the $m$-mers decrease monotonically for all MLIPs as more and more sub-clusters are augmented to the dataset, as expected. For the 8-atom structures, SOAP-BPNN and MACE exhibits a compromise in the accuracy of the full configurations with addition of the sub-cluster $m$-mers in the training dataset. The RMSEs generally increase as more $m$-mers are incorporated, with MACE exhibiting a more pronounced increase—by a factor of 3.4 in both the energy and force RMSEs upon full augmentation. In PET, the RMSEs eventually decrease to values that are lower than those before any $m$-mers are introduced to the dataset. This showcases the exclusive capability of PET, among the considered MLIPs, to learn the reference body-ordered energetics and further improve the accuracy on the full structures.
@@ VI Body-ordered interpretation of MLIP learning dynamics
Having established that the MLIPs infer their own body-order trends, we now investigate how these trends depend on the composition of the training dataset. We first train the models on the sub-sampled versions of the original training set from Sec. IV in order to probe the learning dynamics of the MLIPs in the context of body-orders. We use dataset proportions ranging from 0.01 to 0.4, and keep the validation and test sets fixed. The learning curves are shared in Figure S2. We also repeat the exercise for datasets exclusively composed of either 10,000 low $\rho$ 8-mers or 10,000 high $\rho$ 8-mers. The resulting model accuracies for the latter two cases are shared in Table S2 and S3.
@@
Most notably, we observe the tendency of MACE to prioritize the use of lower body-orders in its learning, which is also faintly present in SOAP-BPNN, and absent in PET. Similar behaviors persist in when the models are exclusively trained on the low $\rho$ 8-mers or the high $\rho$ 8-mers (see Figures S3-S6). MACE consistently prefers a fast-converging body-order trend, while SOAP-BPNN exhibits a converging trend for the low-density-only case and does not for the high-density-only case, and PET continues to show entirely arbitrary trends, especially in $\tilde{V}^{(m)}_{A}$. We attribute this distinct trend of MACE to the over-representation of the lower body-orders (see Appendix D), which encourages the model to prioritize the use of lower body-ordered in formation in the learning task. We conjecture that the over-representation of lower body-orders restricts the use of the higher body-ordered features or descriptors to its full capacity, resulting in the limited capability of MACE to optimize its body-orders as the dataset is further expanded.
@@ VII Extrapolative performance of MLIPs
One common interpretation of the MBE is that once sufficient convergence is reached in the body-ordered energetics, the expansion can be generalized to any applicable system with good accuracy. To assess this, we evaluate the extrapolative performance of the MLIPs, which have shown varying trends of convergence in their effective body-ordered energetics with respect to $m$. The energy and force RMSEs are computed for a dataset of out-of-distribution 8-mers that have been further sampled from the intermediate density regime of the Cheng et al. [23] bulk hydrogen dataset, between the two density extrema from which the original 8-mers were sampled. The intermediate density 8-mers are organized into quintiles, where the their original bulk hydrogen configurations exhibit densities of 0.720, 0.876, 0.999, 1.14, and 1.25 g/cm3, respectively. Results are presented in Figure 6.
@@
When the models are trained on both low $\rho$ and high $\rho$ 8-mers, errors below 0.1 eV/atom are also observed in all three MLIPs for the out-of-distribution intermediate density 8-mers. The lowest RMSEs are consistently observed for PET, which achieves average RMSEs of 0.0061 eV/atom for the energies and 0.148 eV/Å for the forces across all intermediate density quintiles. Next best performance is observed for MACE, with an energy RMSE of 22.8 meV/atom for the first quintile, below 20 meV/atom for all other quintiles, and an average force RMSE of 0.239 eV/Å across all quintiles. SOAP-BPNN exhibits average RMSEs of 45.4 meV/atom and 0.674 eV/Å for the energies and forces, respectively, which are still well within an order of magnitude from the RMSEs observed for the original test set.
@@
When the models are trained on the low $\rho$ 8-mers only, the out-of-distribution performance becomes worse for all three MLIPs. In fact, an overall increasing trend in the RMSEs is observed from the first intermediate density quintile (closest to low $\rho$) all the way to the high $\rho$ 8-mers. For MACE, RMSEs for the fifth quintile and high $\rho$ are notably worse compared to the other quintiles, exhibiting up to two orders of magnitude difference compared to the low $\rho$ 8-mers. For PET, the RMSEs are relatively more constant across the intermediate density quintiles and high $\rho$, with average values of 29.6 meV/atom and 0.409 eV/Å for the energies and forces. The analogous reverse trend is observed when the models are trained on the high $\rho$ 8-mers only, where the RMSEs for the intermediate $\rho$ and low $\rho$ 8-mers become significantly higher than before, and show an increasing trend from the fifth quintile (closest to high $\rho$) to the first quintile and low $\rho$.
@@
Figure 6 also presents the RMSEs for the intermediate $\rho$ 8-mers when the body-order trends of the models are explicitly resolved to the DFT reference (Sec. V). MACE exhibits performance degradation in all cases, and the degradation is far more pronounced when trained on the low or high $\rho$ 8-mers only. SOAP-BPNN shows reduced RMSEs for the case of high $\rho$ 8-mers only, and slight degradation in the other two cases. PET consistently displays slightly lower RMSEs for the out-of-distribution 8-mers when trained on the body-order-resolved datasets. This showcases the flexibility of PET to learn simultaneously the energetics of target 8-mers as well as their sub-clusters, then use the extra information from the sub-clusters to achieve further improvements in the RMSEs.
@@ VIII Conclusion and outlook
In this study, we have carefully analyzed the behavior of three different MLIPs in terms of their “body-orderedness”, examining the body-ordered energy and force trends for hydrogen clusters extracted from bulk simulations at different densities, comparing between DFT and the MLIPs trained on a number of different dataset makeups. In the reference DFT calculations, as seen in many other systems, we first observed that the MBE of energy shows a non-converging, oscillatory behavior for both “molecular” (low density) and “atomic” (high density) hydrogen clusters. The effect cannot simply be dismissed as an artifact of DFT, as the trend is also reproduced in state-of-the-art FCI calculations. Even though the oscillatory behavior can be explained in terms of the choice of the isolated atom energy as the baseline, higher body-order terms are large also for force-based metrics, which are insensitive to the choice of the baseline. An analysis of the electronic structure of the fragments points to the strong spin correlations as the origin of high body-order terms.
@@
When trained exclusively on 8-mers extracted from realistic bulk structures, the MLIPs all learn an effective MBE that is far from the reference—without any adverse effect on their in-set accuracy. While the effective MBE is largely arbitrary for all models, MACE tends to prioritize the use of lower body-orders for a fast-converging trend. When the sub-clusters are incorporated into the training set to explicitly resolve the body-orders to the reference, MACE and PET quickly converge to the reference body-ordered energetics, but the relatively low descriptive power of SOAP-BPNN limits its accuracy on the $m$-mers and hence its ability to learn the reference MBE. We also observe that explicit body-order resolution degrades the accuracy on the full structures of interest for MACE, while it does not for PET. Contrary to what one might expect, the fast decay of the effective MBE does not translate into more robust extrapolative behavior. Explicitly resolving the body-orders does not improve the extrapolative performance for SOAP-BPNN, degrades it for MACE, and improves it slightly for PET.
@@
While our experiments focus on one, comparatively simple, system, they suggest that there is little value in targeting explicitly the MBE, or in designing models that implicitly favor learning a fast-decaying effective body-ordered decomposition. On the contrary, it appears that an unconstrained architecture such as PET, that does not build upon a hierarchical expansion of the neighbor density correlations but simply aims to achieve a highly expressive approximation of the target, demonstrates consistently the best performance, both for in- and out-of-distribution tasks.
@@
The “paradox” of the MBE is resolved by recognizing that models trained on reasonably stable structures don’t have to reproduce the MBE of the target. The large deviations between the true and empirical trends is a simple consequence of the fact that, for a model trained on those reasonable structures, the fragments that appear in the decomposition are highly distorted, and amount to extrapolative predictions. We speculate that the tendency of MACE to privilege a fast-decaying effective body-ordered energy decomposition may be a consequence of the “contamination” of high-order correlations with low-body-order components. Given that a strategy to “purify” the body-ordered components of the closely related atomic cluster expansion [27], it could be interesting to investigate the behavior of a “purified MACE” architecture, to verify this hypothesis and observe if there are any consequences for accuracy and transferability. Overall, our observations suggest that, despite being an attractive approximation framework, and despite the strong mathematical connection to many widely used MLIP frameworks, the body-order decomposition is not especially useful as a guiding principle to design MLIPs.
@@ Acknowledgments
The calculation inputs, model training inputs, datasets and analysis scripts and notebooks will be made available upon publication of this work.
@@ Appendix A Linear body-ordered models
In linear and kernel-based MLIPs with the locality ansatz, the descriptors are often built to describe atom-centered body-order correlations. One example is Gaussian Approximation Potential (GAP) [14] model utilizing SOAP descriptors, [33] which is a 3-body descriptor. In some special cases (e.g. moment-tensor potential [15], ACE [16]), $\mathbf{x}_{i}$ of Eq. (5) can be expressed in terms of multiple body-ordered components:
@@
In such models, body-ordered descriptors of Eq. (9) often contain self-interacting terms, i.e. contributions where $j_{1}=j_{2}$. In ACE, allowing for such self-correlations, which is sometimes referred to as the “density trick”, is what guarantees favorable scaling with the number of atoms in the environments. Chong et al. [26] have shown that the presence of such self-interactions leads to a model learning behavior where the apparent correspondence between the summands of Eqs. (1) and (6) cannot be captured by the model. More recently, Ho et al. [27] proposed a purification operator that removes the self-correlation contributions from the ACE descriptors, which allows the model to recover the above-mentioned correspondence in some specific cases (see Section 5.1 of Chong et al. [26]).
@@ Appendix C Dependence of body-order convergence on baseline energy
These considerations justify the oscillatory behavior observed for the body-ordered energies, and indicate that adjusting $E_{1}$ to an effective value that zeroes out $\epsilon$ would facilitate the convergence of the expansion—consistent with the derivation in Ref. 64 that assumes the definition of a system-specific, effective energy reference.
@@ Appendix D Over-representation of low body-orders in MACE
Here, we discuss two “channels” that lead to the over-representation of the low body-order contributions in MACE. First, consider the initial features of MACE, which are defined in terms of ACE based on the density trick, as discussed in Section II.1. Since this implies the presence of self-interaction terms in the formalism [16, 21], all higher body-order features also contain effectively lower body-order contributions, which makes the descriptors deviate away from the strictly canonical expression of the body orders. That is, in the summation of Eq. (9), there exist terms where $j_{k}=j_{l}$. Recognizing their presence, we can re-express the descriptors by decomposing their contributions, as done in the following for the example of $\nu=2$:
@@
while $x^{2}_{i,\text{pure}}$ contains all the remaining pure $\nu=2$ terms, such that $j_{1}\neq j_{2}$. Because the summation in $x^{2}_{i,\text{self}}$ runs over only one indexes, the self interaction term is a contribution with an effective lower body-order. Extending this argument to every increasing $\nu$, this proves that the basic ACE features, grounded in the density trick, implies that every $\nu$-term contains all the lower body-orders. As such, also every feature in MACE have contains an over-representation of lower orders. Moreover, as shown in Ref. [65] (see Table 2), expansion terms containing self-interaction give rise to ill-conditioned representation. Therefore, one can infer that the terms for lower body order terms are of magnitude which is comparable with the one of the pure ones.
The second channel is provided by the update function, containing also a residual connection, used in the MACE architecture as defined in Ref. [37], which both contribute to the over-representation of the lower body-orders in the effective descriptors. The update function is defined as (from Ref. [37], Eq. (12)):

### v2

@@ V Explicit resolution of reference body-orders
Figure 3 shows that the DFT body-order trends are quickly captured by both MACE and PET as the $m$-mers are added to the training set. In fact, for both $\rho$ 8-mers, near-complete resolution is achieved for $m\leq 6$ at 0.01 added $m$-mer proportion, and further increase in the proportion quickly resolves the $6\leq m\leq 8$ contributions, faster for PET than MACE. This proves that while the MACE and PET architectures have sufficient flexibility to learn the reference body-order trends, without explicit resolution of the body-orders, their inherent tendencies result in trends that are drastically different from the reference. In SOAP-BPNN, $m\leq 4$ contributions are also resolved quickly at 0.01 added $m$-mer proportion, akin to the other two models. SOAP-BPNN, however, struggles to capture the higher body-order contributions at all added $m$-mer proportions, although it still gradually approaches the reference values. Better resolution is achieved for high $\rho$ 8-mers than the low $\rho$ 8-mers. In all MLIPs, the resolution takes place consistently faster for lower $m$, which must be related to how the dimensionality of the potential energy surfaces of $m$-mers exponentially increases with $m$.
Figure 4 shows the energy and force RMSEs for the 8-mers and their fragments. The RMSEs for the $m$-mers decrease monotonically for all MLIPs as more and more sub-clusters are augmented to the dataset, as expected. For the 8-atom structures, SOAP-BPNN and MACE exhibits a compromise in the accuracy of the full configurations with addition of the sub-cluster $m$-mers in the training dataset. The RMSEs generally increase as more $m$-mers are incorporated, with MACE exhibiting a more pronounced increase—by a factor of 3.4 in both the energy and force RMSEs upon full augmentation. In PET, the RMSEs eventually decrease to values that are lower than those before any $m$-mers are introduced to the dataset. Among the considered MLIPs, PET is the only architecture for which learning the reference body-ordered energetics further improves the accuracy on the full structures.
@@ VI Body-ordered interpretation of MLIP learning dynamics
Having established that in the absence of explicit fragments in the training set, the MLIPs infer their own body-order trends, we now investigate how these trends depend on the composition of the training dataset. We first train the models on the sub-sampled versions of the original training set from Sec. IV in order to probe the learning dynamics of the MLIPs in the context of body-orders. We use dataset proportions ranging from 0.01 to 0.4, and keep the validation and test sets fixed. The learning curves are shared in Figure S3. We also repeat the exercise for datasets exclusively composed of either 10,000 low $\rho$ 8-mers or 10,000 high $\rho$ 8-mers. The resulting model accuracies for the latter two cases are shared in Table S2 and S3.
@@
Most notably, we observe the tendency of MACE to prioritize the use of lower body-orders in its learning, which is also faintly present in SOAP-BPNN, and absent in PET. Similar behaviors persist in when the models are exclusively trained on the low $\rho$ 8-mers or the high $\rho$ 8-mers (see Figures S4-S7). MACE consistently prefers a fast-converging body-order trend, while SOAP-BPNN exhibits a converging trend for the low-density-only case and does not for the high-density-only case, and PET continues to show entirely arbitrary trends, especially in $\tilde{V}^{(m)}_{A}$. We attribute this distinct trend of MACE to the over-representation of the lower body-orders (see Appendix D), which encourages the model to prioritize the use of lower body-ordered in formation in the learning task. We conjecture that the over-representation of lower body-orders restricts the use of the higher body-ordered features or descriptors to its full capacity, resulting in the limited capability of MACE to optimize its body-orders as the dataset is further expanded.
One may wonder how much these trends depend on the details of an architecture and the choice of hyperparameters. As we show in the Supplementary Material, increasing the number of channels or correlation order $\nu$ of MACE modulates the body-order trend but doesn’t change the overall fast-decaying tendency. Using nonlinear interaction blocks [Batatia2025] in MACE alleviates the model from converging trends and induces an oscillatory, diverging behavior in $\tilde{V}^{(m)}_{A}$ more similar to PET. Changing the token size for PET leads to changes in the body-order terms but not in their qualitative behavior, which remains oscillatory and without a clear relation to that of the underlying DFT reference.
@@ VII Extrapolative performance of MLIPs
One common interpretation of the MBE is that once sufficient convergence is reached in the body-ordered energetics, the expansion can be generalized to any applicable system with good accuracy. To assess this, we evaluate the extrapolative performance of the MLIPs, which have shown varying trends of convergence in their effective body-ordered energetics with respect to $m$. The energy and force RMSEs are computed for a dataset of out-of-distribution 8-mers that have been further sampled from the intermediate density regime of the Cheng et al. [Cheng2020] bulk hydrogen dataset, between the two density extrema from which the original 8-mers were sampled. The intermediate density 8-mers are organized into quintiles, where the their original bulk hydrogen configurations exhibit densities of 0.720, 0.876, 0.999, 1.14, and 1.25 g/cm3, respectively. Results are presented in Figure 6.
@@
When the models are trained on both low $\rho$ and high $\rho$ 8-mers, errors below 0.1 eV/atom are also observed in all three MLIPs for the out-of-distribution intermediate density 8-mers. The lowest RMSEs are consistently observed for PET, which achieves average RMSEs of 0.0061 eV/atom for the energies and 0.148 eV/Å for the forces across all intermediate density quintiles. Next best performance is observed for MACE, with an energy RMSE of 0.0228 eV/atom for the first quintile, below 0.02 eV/atom for all other quintiles, and an average force RMSE of 0.239 eV/Å across all quintiles. SOAP-BPNN exhibits average RMSEs of 0.0454 eV/atom and 0.674 eV/Å for the energies and forces, respectively, which are still well within an order of magnitude from the RMSEs observed for the original test set.
@@
When the models are trained on the low $\rho$ 8-mers only, the out-of-distribution performance becomes worse for all three MLIPs. In fact, an overall increasing trend in the RMSEs is observed from the first intermediate density quintile (closest to low $\rho$) all the way to the high $\rho$ 8-mers. For MACE, RMSEs for the fifth quintile and high $\rho$ are much worse than the other quintiles, exhibiting up to two orders of magnitude difference compared to the low $\rho$ 8-mers. For PET, the RMSEs are relatively more constant across the intermediate density quintiles and high $\rho$, with average values of 0.0296 eV/atom and 0.409 eV/Å for the energies and forces. The analogous reverse trend is observed when the models are trained on the high $\rho$ 8-mers only, where the RMSEs for the intermediate $\rho$ and low $\rho$ 8-mers become significantly higher than before, and show an increasing trend from the fifth quintile (closest to high $\rho$) to the first quintile and low $\rho$.
@@
Figure 6 also presents the RMSEs for the intermediate $\rho$ 8-mers when the body-order trends of the models are explicitly resolved to the DFT reference (Sec. V). MACE exhibits performance degradation in all cases, and the degradation is far more pronounced when trained on the low or high $\rho$ 8-mers only. SOAP-BPNN shows reduced RMSEs for the case of high $\rho$ 8-mers only, and slight degradation in the other two cases. PET consistently displays slightly lower RMSEs for the out-of-distribution 8-mers when trained on the body-order-resolved datasets. This showcases the flexibility of PET to learn simultaneously the energetics of target 8-mers as well as their sub-clusters, then use the extra information from the sub-clusters to achieve further improvements in the RMSEs. In the Supplementary Material, we show that these trends are generally robust to changes in the hyperparameters of MACE and PET, except when nonlinear interaction blocks are used in MACE, which mitigates the previously observed performance degradation with body-order resolution to the DFT reference.
@@ VIII Conclusion and outlook
In this study, we have carefully analyzed the behavior of three different MLIPs in terms of their “body-orderedness”, examining the body-ordered energy and force trends for hydrogen clusters extracted from bulk simulations at different densities, comparing between DFT and the MLIPs trained on a number of different dataset makeups. In the reference DFT calculations, as seen in many other systems, we first observed that the MBE of energy shows a non-converging, oscillatory behavior for both “molecular” (low density) and “atomic” (high density) hydrogen clusters. The effect cannot simply be dismissed as an artifact of DFT, as the trend is also reproduced in state-of-the-art DMRG calculations. Even though the oscillatory behavior can be explained in terms of the choice of the isolated atom energy as the baseline, higher body-order terms are large also for force-based metrics, which are insensitive to the choice of the baseline. An analysis of the electronic structure of the fragments points to the strong spin correlations as the origin of high body-order terms.
@@
When trained exclusively on 8-mers extracted from realistic bulk structures, the MLIPs all learn an effective MBE that is far from the reference—without any adverse effect on their in-domain accuracy. While the effective MBE is largely arbitrary for all models, MACE tends to prioritize the use of lower body-orders for a fast-converging trend. When the sub-clusters are incorporated into the training set to explicitly resolve the body-orders to the reference, MACE and PET quickly converge to the reference body-ordered energetics, but the relatively low descriptive power of SOAP-BPNN limits its accuracy on the $m$-mers and hence its ability to learn the reference MBE. We also observe that explicit body-order resolution degrades the accuracy on the full structures of interest for MACE, while it does not for PET. Contrary to what one might expect, the fast decay of the effective MBE does not translate into more robust extrapolative behavior. Explicitly resolving the body-orders does not improve the extrapolative performance for SOAP-BPNN, degrades it for MACE, and improves it slightly for PET.
@@
While our experiments focus on one comparatively simple system, they suggest that there is little value in targeting explicitly the MBE, or in designing models that implicitly favor learning a fast-decaying effective body-ordered decomposition. On the contrary, it appears that an unconstrained architecture such as PET, that does not build upon a hierarchical expansion of the neighbor density correlations but simply aims to achieve a highly expressive approximation of the target, demonstrates consistently the best performance, both for in- and out-of-distribution tasks.
@@
The “paradox” of the MBE is resolved by recognizing that models trained on reasonably stable structures don’t have to reproduce the MBE of the target. The large deviations between the true and empirical trends is a simple consequence of the fact that, for a model trained on those reasonable structures, the fragments that appear in the decomposition are highly distorted and amount to extrapolative predictions. We also speculate that the tendency of MACE to privilege a fast-decaying effective body-ordered energy decomposition may be a consequence of the “contamination” of high-order correlations with low-body-order components. We have shown in the Supplementary Material that the newly proposed nonlinear interaction blocks of MACE can alleviate this effect to an extent. [Batatia2025] Given the strategy to “purify” the body-ordered components of the closely related atomic cluster expansion [Ho2024], it could be interesting to investigate the behavior of a “purified MACE” architecture, to verify our hypothesis and observe if there are any consequences for accuracy and transferability. Overall, our observations suggest that, despite being an attractive approximation framework, and despite the strong mathematical connection to many widely used MLIP frameworks, the body-order decomposition is not especially useful as a guiding principle to design MLIPs.
Supplementary Material
A supplementary PDF document is provided, which contains further details of the ab initio calculations, RMSE values, and results of additional experiments with variations in the dataset, baseline energy $E_{1}$, and ML model hyperparameters.
@@ Acknowledgments
The model training inputs, datasets and analysis scripts and notebooks are provided in a Materials Cloud [Talirz2020] repository under the following DOI: 10.24435/materialscloud:q7-da.
@@ Appendix A Linear body-ordered models
In linear and kernel-based MLIPs with the locality ansatz, the descriptors are often built to describe atom-centered body-order correlations. One example is Gaussian Approximation Potential (GAP) [Bartok2010] model utilizing SOAP descriptors, [Bartok2013] which is a 3-body descriptor. In some special cases (e.g. moment-tensor potential [Shapeev2016], ACE [drau19prb]), $\mathbf{x}_{i}$ of Eq. (5) can be expressed in terms of multiple body-ordered components:
@@
In such models, body-ordered descriptors of Eq. (9) often contain self-interacting terms, i.e. contributions where $j_{1}=j_{2}$. In ACE, allowing for such self-correlations, which is sometimes referred to as the “density trick”, is what guarantees favorable scaling with the number of atoms in the environments. Chong et al. [Chong2025] have shown that the presence of such self-interactions leads to a model learning behavior where the apparent correspondence between the summands of Eqs. (1) and (6) cannot be captured by the model. More recently, Ho et al. [Ho2024] proposed a purification operator that removes the self-correlation contributions from the ACE descriptors, which allows the model to recover the above-mentioned correspondence in some specific cases (see Section 5.1 of Chong et al. [Chong2025]).
@@ Appendix C Dependence of body-order convergence on baseline energy
These considerations justify the oscillatory behavior observed for the body-ordered energies, and indicate that adjusting $E_{1}$ to an effective value that zeroes out $\epsilon$ would facilitate the convergence of the expansion—consistent with the derivation in Ref. Thomas2022 that assumes the definition of a system-specific, effective energy reference.
@@ Appendix D Over-representation of low body-orders in MACE
Here, we discuss two “channels” that lead to the over-representation of the low body-order contributions in MACE. First, consider the initial features of MACE, which are defined in terms of ACE based on the density trick, as discussed in Section II.1. Since this implies the presence of self-interaction terms in the formalism [drau19prb, musi+21cr], all higher body-order features also contain effectively lower body-order contributions, which makes the descriptors deviate away from the strictly canonical expression of the body orders. That is, in the summation of Eq. (9), there exist terms where $j_{k}=j_{l}$. Recognizing their presence, we can re-express the descriptors by decomposing their contributions, as done in the following for the example of $\nu=2$:
@@
while $x^{2}_{i,\text{pure}}$ contains all the remaining pure $\nu=2$ terms, such that $j_{1}\neq j_{2}$. Because the summation in $x^{2}_{i,\text{self}}$ runs over only one indexes, the self interaction term is a contribution with an effective lower body-order. Extending this argument to every increasing $\nu$, this proves that the basic ACE features, grounded in the density trick, implies that every $\nu$-term contains all the lower body-orders. As such, also every feature in MACE have contains an over-representation of lower orders. Moreover, as shown in Ref. [duss+22jcp] (see Table 2), expansion terms containing self-interaction give rise to ill-conditioned representation. Therefore, one can infer that the terms for lower body order terms are of magnitude which is comparable with the one of the pure ones.
The second channel is provided by the update function, containing also a residual connection, used in the MACE architecture as defined in Ref. [bata+22nips], which both contribute to the over-representation of the lower body-orders in the effective descriptors. The update function is defined as (from Ref. [bata+22nips], Eq. (12)):